# Adversarial Paraphrasing: A Universal Attack for Humanizing AI-Generated Text

### v1

@@ 1 Introduction
Figure 1: An overview of our universal and training-free framework for humanizing AI text. At every auto-regressive step of adversarial paraphrasing, we search for the most human-like text tokens from the set of top candidate tokens sampled by the paraphraser LLM. We use the guidance signal from an AI text detector to select the most human-like text tokens. The token generation iterations continue until the paraphrasing is finished. (i.e. [EOS] token is sampled)
@@
In this paper, we introduce a universal, training-free framework for humanizing AI-generated text, a novel attack designed to effectively and efficiently rephrase AI-generated text into “human-like text” that evades a wide range of detectors with greater success when compared to the existing evasion attack methods. Our approach, Adversarial Paraphrasing, leverages an existing instruction fine-tuned LLM (LLaMA-3-8B configured as a paraphraser via a custom system prompt) not just for paraphrasing, but for adversarially humanizing the text under the guidance of a trained AI text detector (see Figure 1). During the paraphrasing process, at each token generation step, we consider the top likely next tokens proposed by the paraphrasing LLM. Then, instead of standard decoding, we score each of these potential continuations using a guidance AI text detector. We then select the token that leads the sequence to be classified as most human-like by the detector, akin to a targeted, detector-guided beam search at depth one. Some example results are shown in Table 1. This method operationalizes the goal of humanization directly within the generation process, falling under the umbrella of controlled text generation (zhou2023controlled, ; liang2024controllable, ; dathathri2019plug, ), where the desired attribute is “human-likeness” as judged by a trained detector.
We demonstrate that our adversarial paraphrasing strategy is not only more effective than baseline methods such as simple or recursive paraphrasing, but unlike other methods, it is also universally transferable to other diverse detection methods, including neural network-based, zero-shot, and watermarking-based detectors. This justifies the universality of our framework. For example, while simple paraphrasing ironically increases true positive at 1% false positive (T@1%F) by 8.57% on RADAR hu2023radar and 15.03% on Fast-DetectGPT bao2023fast , our adversarial approach, guided by OpenAI-RoBERTa-Large solaiman2019release , reduces T@1%F by 64.49% and a striking 98.96%, respectively. Evaluated across eight detectors from three different categories, our method achieves an average T@1%F reduction of 87.88%, all while maintaining minimal degradation in text quality. We comprehensively analyze the tradeoff between our attack and the text quality using multiple automated studies.
@@ 3 Adversarial Paraphrasing for Universal Humanization of AI Text
0:   Input: System instruction $sys$, AI-generated text $x_{:n}$, top-p and top-k token masking methods $top\_p$ and $top\_k$
@@
7:     for $k=1$ to length($candidates$do
8:        $scores$.append($\mathcal{D}(y_{:m}\oplus candidates[k])$)
@@
In this section, we present our adversarial paraphrasing framework, designed to universally paraphrase AI-generated text into human-like text that evades detection on various detectors. Algorithm 1 outlines our method, and Figure 1 provides an illustrative visual overview of each step in the iterative paraphrasing process. Our approach auto-regressively generates paraphrased text using a paraphrasing model $\mathcal{P}:\mathcal{X}\to\mathcal{X}$, guided by a neural network-based detector $\mathcal{D}:\mathcal{X}\to[0,1]$, where $\mathcal{X}$ denotes the space of natural language texts. The model $\mathcal{P}$ outputs the next token logit distribution $p(\cdot|x)\in\mathbb{R}^{d}$ for any input $x\in\mathcal{X}$, where $d$ represents the vocabulary size. In the standard setting, the paraphrasing model would multinomially sample the next token from this distribution. In our method, however, token selection is influenced by the detector $\mathcal{D}$, which assigns lower scores (i.e., closer to 0) to more human-like texts.
Overview of the Algorithm. As shown in line 1 of Algorithm 1, we initialize the output string as empty, i.e., $y=\text{``''}$ “”. The algorithm then enters an auto-regressive loop (lines 2–15), continuing until the end-of-sentence ([EOS]) token is generated. At each iteration, the paraphraser computes the logit distribution for the next token (line 3). To narrow down the candidate set, we apply top-$p$ filtering to only select the top tokens whose cumulative probability exceeds a certain threshold $p$, and top-$k$ filtering to cap the maximum number of candidate tokens (line 4). In line 5, we decode the filtered logits into their corresponding textual representations using the decoding function $\mathcal{T}$. From lines 6 to 9, we score each candidate by appending it to the current output and evaluating the resulting text using the detector $\mathcal{D}$. The token associated with the lowest detector score—i.e., the most human-like continuation—is selected and appended to the output sequence (lines 10 and 14). The loop continues until the [EOS] token is generated.
@@
The key intuition is that most, if not all, high-performing detectors tend to converge toward a common distribution that characterizes human-authored text, in an effort to minimize false positives. Consequently, if a paraphraser is guided to evade detection by a well-trained detector, its outputs may naturally align more closely with this shared human-like distribution. As a result, the generated text becomes more difficult to detect not only by the detector used for guidance, but also by other detectors—since they are all ideally calibrated to the same underlying distribution of human-written text. This property makes our adversarial paraphrases broadly transferable across different detectors.
@@ 4.1 Setup
Datasets. For non-watermark-based detectors, we use MAGE li2023mage as our primary evaluation dataset due to its rich diversity of text sources. We randomly sample 2000 AI-generated texts and 2000 human-written texts from MAGE while ensuring that each text is $\sim$100 to 500 tokens in length. For watermark-based detectors, we construct “watermarked” datasets using a watermarked LLaMA-3.1-8B-Instruct grattafiori2024llama3herdmodels . Specifically, we input the model with the first 20 words of each of the 2000 AI texts as prefix, and let it generate watermarked text $\sim$200 to 600 tokens in length. We report detailed token statistics for all datasets used in Appendix B.
@@
{tblr}
width=rowsep=0.5pt, colspec=X[c] c, column1 = halign=justify, column2 = halign=c, row1 = halign=c, hline1,11 = -0.08em, hline2,5,8 = -, Text & Rating
Original AI Text. There doesn’t seem to be a whole lot of information available about DOCSIS 3.1 yet, but if my prior experience can lend a hand towards this question, I would venture The DOCSIS 3.1 issue has a number of things to keep in mind. First, the DOCSIS 3.1 expansion is only available for initial support, and this upgrade is apparently working on [continues…] –
@@
Original AI Text. No, addiction is much more than simply a habit; it is a chronic and progressive medical disorder. While habits can be formed through repeated use of drugs or alcohol, addiction is a physical and psychological response to those drugs that causes compulsive drug seeking and use despite the negative consequences. It has been suggested that addiction is [continues…] –
@@
Original AI Text. The occurrence of near misses was linked to socio-economic and clinical factors as well as issues with care, highlighting the critical need for measures to enhance these indicators. Furthermore, we recommend a more effective incorporation of this topic into the medical course curriculum to emphasize the significance of preventing near misses through appropriate medical training [continues…] –
@@ 5 Quality Evaluation of Paraphrased Texts
We conduct a comprehensive evaluation to investigate the impact of adversarial paraphrasing on the perceived quality of AI-generated text. For this, we randomly sample 100 texts each from our three datasets (MAGE, KGW watermarked MAGE, and Unigram watermarked MAGE) and analyze them with three complementary studies: (1) perplexity scores (PPL), (2) auto-rated quality from GPT-4o comparing paraphrased texts to their original AI counterparts, and (3) head-to-head win rates, also assessed by GPT-4o, comparing adversarial paraphrasing against simple paraphrasing. We provide representative examples of paraphrases in Table 1, with a much broader set of examples included in Appendix C to support qualitative manual inspection. Our findings highlight a nuanced trade-off between evading detectors and preserving textual quality.
@@
Perplexity Analysis. We assess perplexity using LLaMA-3.1–8B-Instruct grattafiori2024llama3herdmodels , comparing the original AI-generated text, simple paraphrases, and adversarial paraphrases. The results are summarized in Table 3. As shown in the table, human-written text typically exhibits higher perplexity than AI text, as human language tends to deviate more from the statistical regularities learned by LLMs. After applying simple paraphrasing, we observe a substantial improvement in the perplexity of AI text. This may be attributed to the fact that the model used for paraphrasing (LLaMA-3.1) is superior to the LLMs used for generating AI texts in the MAGE dataset (e.g. LLaMA). In contrast, adversarial paraphrasing yields perplexity scores that are comparable to the human texts from MAGE, which is reasonable given that our objective is to humanize AI texts.
@@
Auto-Rating with GPT-4o. In order to evaluate the quality of paraphrases, we employ GPT-4o openai2024gpt4technicalreport , as a judge LLM for automatic quality evaluations vu2024foundational ; fu2023gptscore ; kim2024prometheus with custom system and user prompts (see Appendix D). The judge model is tasked to rate the paraphrases when compared to their original corresponding AI text on a Likert scale of 1-5, in terms of quality and semantic similarity. The first row of Figure 5 shows the quality ratings for both baseline simple paraphrasing and adversarial paraphrasing. Though we observe a slight tradeoff in the text quality when compared to simple paraphrasing, in 87% of the times—averaged across all three datasets and four guidance detectors—adversarial paraphrases were rated 4 or 5 out of 5 (see Appendix E for detailed rating for each guidance detector and dataset). Note that the error bar in the figure shows that this difference between simple and adversarial paraphrasing is not statistically significant. While adversarial paraphrasing can lead to a higher perplexity score when compared to simple paraphrasing, our auto-rater study shows that both the paraphrases have a comparable text quality, making our attack a practical one.
@@ Acknowledgment
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, an award from meta 314593-00001 and an award from Capital One.
@@ Appendix A Comparison with Watermark Stealing
Appendix B Detailed Token Statistics for Evaluated Datasets

### v2

@@ 1 Introduction
Figure 1: An overview of our universal and training-free framework for humanizing AI text. At every auto-regressive step of adversarial paraphrasing, using the guidance from an AI text detector, we search for the token with lowest “AI-score” from the set of top candidate tokens sampled by the paraphraser LLM. The token generation iterations continue until the paraphrasing is finished. (i.e. [EOS] token is sampled)
@@
In this paper, we introduce a universal, training-free framework for humanizing AI-generated text, a novel attack designed to effectively and efficiently rephrase AI-generated text to evade a wide range of detectors with greater success when compared to the existing attack evasion methods. Our approach, Adversarial Paraphrasing, leverages an existing instruction fine-tuned LLM (LLaMA-3-8B configured as a paraphraser via a custom system prompt) not just for paraphrasing, but for adversarially humanizing the text under the guidance of a trained AI text detector (see Figure 1). During the paraphrasing process, at each token generation step, we consider the top likely next tokens proposed by the paraphrasing LLM. Then, instead of standard decoding, we score each of these potential continuations using a guidance AI text detector. We then select the token that leads the sequence to be predicted with the lowest “AI score” (or equivalently, the highest “human score”) by the detector, akin to a targeted, detector-guided beam search at depth one. Some example results are shown in Table 1. This method operationalizes the goal of humanization directly within the generation process, falling under the umbrella of controlled text generation (zhou2023controlled, ; liang2024controllable, ; dathathri2019plug, ).
We demonstrate that our adversarial paraphrasing strategy is not only more effective than baseline methods such as simple or recursive paraphrasing, but unlike other methods, it is also universally transferable to other diverse detection methods, including neural network-based, zero-shot, and watermarking-based detectors. This justifies the universality of our framework. For example, while simple paraphrasing ironically increases true positive rate at 1% false positive rate (T@1%F) by 8.57% on RADAR hu2023radar and 15.03% on Fast-DetectGPT bao2023fast , our adversarial approach, guided by OpenAI-RoBERTa-Large solaiman2019release , reduces T@1%F by 64.49% and a striking 98.96%, respectively. Evaluated across eight detectors from three different categories, our method achieves an average T@1%F reduction of 87.88%, all while maintaining minimal degradation in text quality. We comprehensively analyze the tradeoff between our attack and the text quality using multiple automated studies.
@@ 3 Adversarial Paraphrasing for Universal Humanization of AI Text
0:Input: System instruction syssys, AI-generated text $x_{:n}$, top-p and top-k token masking methods top_ptop\_p and top_ktop\_k
@@
7:  for $k=1$ to length(candidatescandidates) do
[…]
@@
In this section, we present our adversarial paraphrasing framework, designed to universally paraphrase AI-generated text to evade detection on various detectors. Algorithm 1 outlines our method, and Figure 1 provides an illustrative visual overview of each step in the iterative paraphrasing process. Our approach auto-regressively generates paraphrased text using a paraphrasing model $\mathcal{P}:\mathcal{X}\to\mathcal{X}$, guided by a neural network-based detector $\mathcal{D}:\mathcal{X}\to[0,1]$, where $\mathcal{X}$ denotes the space of natural language texts. The model $\mathcal{P}$ outputs the next token logit distribution $p(\cdot|x)\in\mathbb{R}^{d}$ for any input $x\in\mathcal{X}$, where $d$ represents the vocabulary size. In the standard setting, the paraphrasing model would multinomially sample the next token from this distribution. In our method, however, token selection is influenced by the detector $\mathcal{D}$, which assigns lower AI-scores (i.e., closer to 0) to more “human-like” texts in the eyes of the detector.
Overview of the Algorithm. As shown in line 1 of Algorithm 1, we initialize the output string as empty, i.e., $y=\text{``''}$. The algorithm then enters an auto-regressive loop (lines 2–15), continuing until the end-of-sentence ([EOS]) token is generated. At each iteration, the paraphraser computes the logit distribution for the next token (line 3). To narrow down the candidate set, we apply top-$p$ filtering to only select the top tokens whose cumulative probability exceeds a certain threshold $p$, and top-$k$ filtering to cap the maximum number of candidate tokens (line 4). In line 5, we decode the filtered logits into their corresponding textual representations using the decoding function $\mathcal{T}$. From lines 6 to 9, we score each candidate by appending it to the current output and evaluating the resulting text using the detector $\mathcal{D}$. The token associated with the lowest detector score is selected and appended to the output sequence (lines 10 and 14). The loop continues until the [EOS] token is generated.
@@
The key intuition is that most, if not all, high-performing detectors tend to converge toward a common distribution that characterizes human-authored text, in an effort to minimize false positives. Consequently, if a paraphraser is guided to evade detection by a well-trained detector, its outputs may naturally align more closely with this shared human text distribution. As a result, the generated text becomes more difficult to detect not only by the detector used for guidance, but also by other detectors—since they are all ideally calibrated to the same underlying distribution of human-written text. This property makes our adversarial paraphrases broadly transferable across different detectors.
@@ 4.1 Setup
Datasets. For non-watermark-based detectors, we use MAGE li2023mage as our primary evaluation dataset due to its rich diversity of text sources. We randomly sample 2000 AI-generated texts and 2000 human-written texts from MAGE while ensuring that each text is $\sim$100 to 500 tokens in length. For watermark-based detectors, we construct “watermarked” datasets using a watermarked LLaMA-3.1-8B-Instruct grattafiori2024llama3herdmodels . Specifically, we input the model with the first 20 words of each of the 2000 AI texts as prefix, and let it generate watermarked text $\sim$200 to 600 tokens in length. We report detailed token statistics for all datasets used in Appendix C.
@@
Text Rating
Original AI Text. There doesn’t seem to be a whole lot of information available about DOCSIS 3.1 yet, but if my prior experience can lend a hand towards this question, I would venture The DOCSIS 3.1 issue has a number of things to keep in mind. First, the DOCSIS 3.1 expansion is only available for initial support, and this upgrade is apparently working on [continues…]
@@
Original AI Text. No, addiction is much more than simply a habit; it is a chronic and progressive medical disorder. While habits can be formed through repeated use of drugs or alcohol, addiction is a physical and psychological response to those drugs that causes compulsive drug seeking and use despite the negative consequences. It has been suggested that addiction is [continues…]
@@
Original AI Text. The occurrence of near misses was linked to socio-economic and clinical factors as well as issues with care, highlighting the critical need for measures to enhance these indicators. Furthermore, we recommend a more effective incorporation of this topic into the medical course curriculum to emphasize the significance of preventing near misses through appropriate medical training [continues…]
@@ 4.2 Effectiveness and Universality of Adversarial Paraphrasing for Humanizing AI Text
4.3 Efficiency of Adversarial Paraphrasing
Compared to simple paraphrasing, our method requires running a surrogate AI text detector at each decoding step by design. This naturally raises the question of how much latency this process introduces and whether it affects efficiency.
To evaluate this, we conducted adversarial paraphrasing on 100 randomly selected text samples, with five trials per configuration, using all four guidance detectors. To simplify measurement and
Method Run time
Simple Paraphrase $7.18\pm 0.13$
AdvPara (roblarge) $10.20\pm 0.18$
AdvPara (robbase) $8.64\pm 0.11$
AdvPara (mage) $16.71\pm 0.74$
AdvPara (radar) $9.69\pm 0.20$
Table 3: Per-sample run time (in seconds) of adversarial paraphrasing over 100 randomly selected samples.
report time on a per-sample basis (rather than per batch), we used a batch size of 1 during these trials. In practical settings, including the main experiments described in Section 4.2, we employ larger batch sizes, which substantially reduce total runtime. The average paraphrasing time per sample across five trials is reported as the mean ± standard deviation (in seconds [^1] Note that runtime per sample may vary with sequence length. To ensure a fair comparison—i.e., that AdvPara generates a similar number of tokens as simple paraphrasing—we report the mean and standard deviation of token counts in Appendix B.) in Table 3. As shown, most guidance detectors introduce only minor latency compared to simple paraphrasing. The higher latency observed with MAGE arises from its LongFormer-based architecture, which requires longer inference time than the RoBERTa-based models.
Overall, latency is primarily determined by the computational complexity of the guidance detector. From a computational cost perspective, the detector adds minimal overhead in terms of FLOPs relative to the paraphrasing LLM, which dominates total computation. In our experiments, the paraphrasing LLM contains approximately 8 billion parameters, whereas the detectors range from 100 to 350 million parameters—less than 5% of the paraphraser’s size for the larger detectors and under 2% for the smaller ones.
4.4 Ablation Studies
Other than the guidance detector choice, the two main hyper-parameters related to our work are the $p$ and $k$ used in top-p and top-k masking. In the above, we have used $p=0.99$ and $k=50$, where $k=50$ is also the default $k$ value set in HuggingFace Transformer library. To show how their settings
Table 4: Attack effectiveness under different values of $p$ and $k$ used in top-p and top-k masking.
affect our attack, we use OpenAI-RoBERTa-Large as the guidance detector and run ablation studies on the value of $p$ and $k$ on 500 randomly selected texts. Defaulting to $k=50$ and $p=0.99$ when varying the other, we report the mean AUC and TPR@1%FPR across all 8 deploy detectors in Table 4.4.
Fixing $k$ at 50, we can observe that when $p$ exceeds 0.99, the marginal gains in attack effectiveness diminish relative to the losses in text quality. Additionally, as one can imagine, further increasing slows down generation speed. Therefore, we select $p=0.99$ as a balanced trade-off among attack effectiveness, text quality, and generation speed. With $p$ fixed at 0.99, the values of $k$ show relatively smaller influence, and mostly lead to comparable results. Hence we just stick with the default value in the HuggingFace Transformer library.
@@ 5 Quality Evaluation of Paraphrased Texts
We conduct a comprehensive evaluation to investigate the impact of adversarial paraphrasing on the perceived quality of AI-generated text, focusing on both semantic equivalence to the original text and clarity, fluency, and naturalness of the text itself. For this, we randomly sample 100 texts each from our three datasets (MAGE, KGW watermarked MAGE, and Unigram watermarked MAGE) and analyze them with four complementary studies: (1) perplexity scores (PPL), (2) SBERT reimers2019sentencebertsentenceembeddingsusing semantic similarity, (3) auto-rated quality from GPT-4o comparing paraphrased texts to their original AI counterparts, and (4) head-to-head win rates, also assessed by GPT-4o, comparing adversarial paraphrasing against simple paraphrasing. We provide representative examples of paraphrases in Table 1, with a much broader set of examples included in Appendix F to support qualitative manual inspection. Our findings highlight a nuanced trade-off between evading detectors and preserving textual quality.
@@
Perplexity Analysis. We assess perplexity using LLaMA-3.1–8B-Instruct grattafiori2024llama3herdmodels , comparing the original AI-generated text, simple paraphrases, and adversarial paraphrases. The results are summarized in Table 5. As shown in the table, human-written text typically exhibits higher perplexity than AI text, as human language tends to deviate more from the statistical regularities
Method SBERT Cos. Sim.
Simple Paraphrase $0.8601\pm 0.0880$
AdvPara (roblarge) $0.8082\pm 0.1006$
AdvPara (robbase) $0.8128\pm 0.0985$
AdvPara (mage) $0.8159\pm 0.0982$
AdvPara (radar) $0.8095\pm 0.1025$
Table 6: Cosine similarity of SBERT embeddings before and after paraphrasing.
learned by LLMs. After applying simple paraphrasing, we observe a substantial improvement in the perplexity of AI text. This may be attributed to the fact that the model used for paraphrasing (LLaMA-3.1) is superior to the LLMs used for generating AI texts in the MAGE dataset (e.g. LLaMA). In contrast, adversarial paraphrasing yields perplexity scores that are comparable to the human texts from MAGE, which is reasonable given that our objective is to humanize AI texts.
SBERT Similarity. A common approach to assessing the semantic equivalence between two texts is to compute the cosine similarity between their SBERT embeddings reimers2019sentencebertsentenceembeddingsusing . Accordingly, we measure the cosine similarity of SBERT embeddings before and after paraphrasing, and compare adversarial paraphrasing with simple paraphrasing. The results are summarized in Table 6. Overall, although there is a slight reduction in the mean cosine similarity for adversarial paraphrasing, the values remain within an acceptable range given the high variance observed across samples.
@@
Auto-Rating with GPT-4o. In order to simulate the human perception of text quality (i.e., alignment with the “gold standard” of fluency and coherence), we employ GPT-4o openai2024gpt4technicalreport , as a judge LLM for automatic quality evaluations vu2024foundational ; fu2023gptscore ; kim2024prometheus with custom system and user prompts (see Appendix G). The judge model is tasked to rate the paraphrases when compared to their original corresponding AI text on a Likert scale of 1-5, in terms of quality and semantic similarity. The first row of Figure 5 shows the quality ratings for both baseline simple paraphrasing and adversarial paraphrasing. Though we observe a slight tradeoff in the text quality when compared to simple paraphrasing, in 87% of the times—averaged across all three datasets and four guidance detectors—adversarial paraphrases were rated 4 or 5 out of 5 (see Appendix H for detailed rating for each guidance detector and dataset). Note that the error bar in the figure shows that this difference between simple and adversarial paraphrasing is not statistically significant. While adversarial paraphrasing can lead to a higher perplexity score when compared to simple paraphrasing, our auto-rater study shows that both the paraphrases have a comparable text quality, making our attack a practical one.
@@ Acknowledgment
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, the ONR PECASE grant N00014-25-1-2378, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, DARPA AIQ grant HR00112590066, and an award from meta 314593-00001.
@@ Appendix A Comparison with Watermark Stealing
Appendix B Text Token Counts in Efficiency Assessment
The exact runtime per sample may vary depending on the sequence length, which is inherently determined by the length of the input text. To ensure a fair comparison—such that adversarial paraphrasing generates a comparable number of tokens to simple paraphrasing—we report the mean and standard deviation of token counts in Table 8. As shown, both simple paraphrasing and adversarial paraphrasing with different guidance detectors produce a similar mean number of tokens. However, adversarial paraphrasing exhibits a larger variance in token counts, reflecting greater variability in output length.
Text Mean Token Count Std Dev
Original texts 173.73 38.46
Simple Paraphrase 170.81 39.69
AdvPara (RoBERTa-Base) 175.30 51.27
AdvPara (RoBERTa-Large) 171.25 45.39
AdvPara (RADAR) 169.68 60.07
AdvPara (MAGE) 164.18 54.39
Table 8: Mean and standard deviation of token counts for the original, simple paraphrased, and adversarially paraphrased texts used in our efficiency evaluation. All token counts were obtained using the LLaMA-3 tokenizer.
Appendix C Detailed Token Statistics for Evaluated Datasets